𝑓(π‘₯) = 2π‘₯^3 + 𝑐π‘₯^2 + 2π‘₯



Suppose 𝑓 is differentiable on ℝ and has two roots. Show that 𝑓′ has at least one root.

1
Expert's answer
2022-04-07T09:43:22-0400

Rewrite the function in the form: "f(x)=x(2x^2+cx+2)". One root of equation "f=0" is "x=0". Thus, equation "2x^2+cx+2=0" has only one root. It means that "c=4" and "2(x+1)^2=0". The root of the latter is "x=-1." Otherwise, we receive two roots or no roots. The derivative of the function "f" is: "f'=6x^2+8x+2". Roots of equation "f'=0" are: "x=\frac{-8\pm4}{12}=-1,-\frac{1}{3}". Thus, there are two roots of equation "f'=0."